National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 1761–1780 of 2520Page 89 of 100
NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1761

National Defence Academy (NDA) 2018• Mathematics• Mathematics

If \({\rm{y}} = {{\rm{e}}^{{{\rm{x}}^2}}}\sin 2{\rm{x}}\), then what is \(\frac{{{\rm{dy}}}}{{{\rm{dx}}}}\) at x = π equal to?

  1. AOption (a)

    \(\left( {1 + {\rm{\pi }}} \right){{\rm{e}}^{{{\rm{\pi }}^2}}}\)

  2. BOption (b)

    \(2{\rm{\pi \;}}{{\rm{e}}^{{{\rm{\pi }}^2}}}\)

  3. COption (c)

    \(2{{\rm{e}}^{{{\rm{\pi }}^2}}}\)

  4. DOption (d)

    \({{\rm{e}}^{{{\rm{\pi }}^2}}}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(2{{\rm{e}}^{{{\rm{\pi }}^2}}}\)

Why this answer is correct

Concept:

Product rule of differentiation:

\(\frac{{\rm{d}}}{{{\rm{dx}}}}\left( {{\rm{f}}\left( {\rm{x}} \right){\rm{g}}\left( {\rm{x}} \right)} \right) = {\rm{f'}}\left( {\rm{x}} \right){\rm{g}}\left( {\rm{x}} \right) + {\rm{f}}\left( {\rm{x}} \right){\rm{g'}}\left( {\rm{x}} \right)\)

Calculation:

Given: \({\rm{y}} = {{\rm{e}}^{{{\rm{x}}^2}}}\sin 2{\rm{x}}\)

To find: \(\frac{{{\rm{dy}}}}{{{\rm{dx}}}}\)

\({\rm{y}} = {{\rm{e}}^{{{\rm{x}}^2}}}\sin 2{\rm{x}}\)

Differentiating both sides w.r.t x

\( \Rightarrow \frac{{{\rm{dy}}}}{{{\rm{dx}}}} = {{\rm{e}}^{{{\rm{x}}^2}}}\left( {2{\rm{x}}} \right)\sin 2{\rm{x}} + \cos 2{\rm{x}} \times \left( 2 \right) \times {{\rm{e}}^{{{\rm{x}}^2}}}{\rm{\;}}\)

\(\Rightarrow {\left( {\frac{{{\rm{dy}}}}{{{\rm{dx}}}}} \right)_{{\rm{x}} = {\rm{\pi }}}} = {{\rm{e}}^{{{\rm{\pi }}^2}}}\left( {2{\rm{\pi }}} \right)\sin \left( {2{\rm{\pi }}} \right) + 2{{\rm{e}}^{{{\rm{\pi }}^2}}}\cos \left( {2{\rm{\pi }}} \right)\)

 \(\Rightarrow {\left( {\frac{{{\rm{dy}}}}{{{\rm{dx}}}}} \right)_{{\rm{x}} = {\rm{\pi }}}} = 2{{\rm{e}}^{{{\rm{\pi }}^2}}}\)

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1762

National Defence Academy (NDA) 2018• Mathematics• Mathematics

What is the solution of (1 +2x) dy – (1 – 2y) dx = 0?

  1. AOption (a)

    x – y – 2xy = c

  2. BOption (b)

    y – x – 2xy = c

  3. COption (c)

    y + x – 2xy = c

  4. DOption (d)

    x + y + 2xy = c

View answer & explanation

Correct answer

Answer: (a)A.

x – y – 2xy = c

Why this answer is correct

Concept:

Formula used:

  • \(\smallint \frac{{{\rm{dx}}}}{{{\rm{a}} + {\rm{bx}}}} = \frac{1}{{\rm{b}}}\ln \left( {{\rm{a}} + {\rm{bx}}} \right) + {\rm{c}}\)

Calculation:

Given: (1 +2x) dy – (1 – 2y) dx = 0

⇒ (1 +2x) dy = (1 – 2y) dx

\(\Rightarrow \frac{{{\rm{dy}}}}{{1 - 2{\rm{y}}}} = \frac{{{\rm{dx}}}}{{1 + 2{\rm{x}}}}\)

On integrating both sides,

\( \Rightarrow \smallint \frac{{{\rm{dy}}}}{{1 - 2{\rm{y}}}} = \smallint \frac{{{\rm{dx}}}}{{1 + 2{\rm{x}}}}{\rm{\;}}\)

\(\Rightarrow - \frac{1}{2}\ln \left( {1 - 2{\rm{y}}} \right) = \frac{1}{2}\ln \left( {1 + 2{\rm{x}}} \right) + {{\rm{c}}_1}\)

⇒ ln (1 + 2x) + ln (1 – 2y) = ln c2

⇒ (1 + 2x) × (1 – 2y) = c2

⇒ 1 – 2y + 2x – 4xy = c2

⇒ 2x – 2y – 4xy = c2 – 1

⇒ x – y – 2xy = c

Where c1 is integration constant and c2 & c are some arbitrary constant

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1763

National Defence Academy (NDA) 2018• Mathematics• Mathematics

What are the order and degree, respectively, of the differential equation \({\left( {\frac{{{{\rm{d}}^3}{\rm{y}}}}{{{\rm{d}}{{\rm{x}}^3}}}} \right)^2} = {{\rm{y}}^4} + {\left( {\frac{{{\rm{dy}}}}{{{\rm{dx}}}}} \right)^5}?\)

  1. AOption (a)

    4, 5

  2. BOption (b)

    2, 3

  3. COption (c)

    3, 2

  4. DOption (d)

    5, 4

View answer & explanation

Correct answer

Answer: (c)C.

3, 2

Why this answer is correct

Concept:

  • Order: The order of a differential equation is the order of the highest derivative appearing in it.
  • Degree: The degree of a differential equation is the power of the highest derivative occurring in it, after the Equation has been expressed in a form free from radicals as far as the derivatives are concerned.

For example, if the equation is

\({\left( {\frac{{{{\rm{d}}^2}{\rm{y}}}}{{{\rm{d}}{{\rm{x}}^2}}}} \right)^3} + \frac{{{\rm{dy}}}}{{{\rm{dx}}}} - {{\rm{x}}^2} = 0\)

Here, order of the highest derivative is 2. And the degree is 3.

Calculation:

Given: \({\left( {\frac{{{{\rm{d}}^3}{\rm{y}}}}{{{\rm{d}}{{\rm{x}}^3}}}} \right)^2} = {{\rm{y}}^4} + {\left( {\frac{{{\rm{dy}}}}{{{\rm{dx}}}}} \right)^5}\)

To find: Order and degree

So, the highest order is \(\frac{{{{\rm{d}}^3}{\rm{y}}}}{{{\rm{d}}{{\rm{x}}^3}}}\)

∴ Order is 3.

Degree of differential equation is the degree highest order. So, degree is 2.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1764

National Defence Academy (NDA) 2018• Mathematics• Mathematics

In a Binomial distribution, the mean is three times its variance. What is the probability of exactly 3 successes out of 5 trials?

  1. AOption (a)

    80/243

  2. BOption (b)

    40/243

  3. COption (c)

    20/243

  4. DOption (d)

    10/243

View answer & explanation

Correct answer

Answer: (a)A.

80/243

Why this answer is correct

Concept:

Binomial distribution: If a random variable X has binomial distribution as B (n, p) with n and p as parameters, then the probability of random variable is given as:

P(X = k) = nCk pk (1 – p)(n – k)

Where, n is number of observations.

p is the probability of success & (1 – p) is probability of failure.

k = 0, 1, 2, …., n

Properties:

  • Mean of the distribution (μX) = n × p
  • The variance (σ2x) = n × p × (1 – p)
  • Standard deviation (σx) = √{np(1 – p)}

Calculation:

Given: Mean (μX) = 3 × variance (σ2x)

n × p = 3 × n × p × (1 – p)

⇒ 1 – p = 1/3

⇒ p = 2/3

Probability of exactly 3 successes = P(X = 3)

= 5C3 p3 (1 – p)(5 – 3)

\(= \frac{{5\; \times \;4\; \times \;3!}}{{2!3!}} \times {\left( {\frac{2}{3}} \right)^3} \times {\left( {\frac{1}{3}} \right)^2}\)

= 10 × 8 × (1/3)5

= 80/243

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1765

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Consider the following

Statements:

  1. P(A̅ ∪ B) = P(A̅) + P (B) – P(A̅ ∩ B)

  2. P(A ∪ B̅ ) = P(B) – P(A ∩ B)

  3. P(A ∩ B) = P(B) P(A|B)

Which of the above statements are correct?

  1. AOption (a)

    1 and 2 only

  2. BOption (b)

    1 and 3 only

  3. COption (c)

    2 and 3 only

  4. DOption (d)

    1, 2, and 3

View answer & explanation

Correct answer

Answer: (b)B.

1 and 3 only

Why this answer is correct

Concept:

Formulas used:

  • P(A ∪ B) = P(A) + P(B) – P(A ∩ B)
  • Conditional Probability:

\({\rm{P}}\left( {{\rm{A|B}}} \right) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}}\)

Calculation:

For first statement,

We know that, P(A ∪ B) = P(A) + P(B) – P(A ∩ B)

So, P(A̅ ∪ B) = P(A̅) + P(B) – P(A̅ ∩ B)

∴ First statement is correct.

For second statement,

We know that, P(A ∪ B) = P(A) + P(B) – P(A ∩ B)

So, P(A ∪ B̅) = P(A) + P(B̅) – P(A ∩ B̅)

⇒ P(A ∪ B̅) = P(A) + 1 – P(B) – {P(A) – P(A ∩ B)

⇒ P(A ∪ B̅) = 1 – P(B) + P(A ∩ B)

∴ Second statement is incorrect.

For third statement,

From conditional probability we know,

\({\rm{P}}\left( {{\rm{A|B}}} \right) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}}\)

⇒ P(A ∩ B) = P(A|B) × P(B)

∴ Third statement is correct.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1766

National Defence Academy (NDA) 2018• Mathematics• Mathematics

If the correlation coefficient between x and y is 0.6 covariance is 27 variance of y is 25, then what is the variance of x?

  1. AOption (a)

    9/5

  2. BOption (b)

    81/25

  3. COption (c)

    9

  4. DOption (d)

    81

View answer & explanation

Correct answer

Answer: (d)D.

81

Why this answer is correct

Concept:

Formulas used:

\({\rm{r}}\left( {{\rm{x}},{\rm{y}}} \right) = \frac{{{\rm{\;Cov}}\left( {{\rm{x}},{\rm{y}}} \right)}}{{{\rm{\sigma }}\left( {\rm{x}} \right){\rm{\sigma }}\left( {\rm{y}} \right)}}
\)

Calculation:

Given:

Correlation coefficient between x and y, r(x, y) = 0.6

Covariance of x and y, Cov(x, y) = 27

Variance of y, Var(y) = σ2(y) = 25

Variance of x, Var(x) = σ2(x)

We know that,\({\rm{r}}\left( {{\rm{x}},{\rm{y}}} \right) = \frac{{{\rm{\;Cov}}\left( {{\rm{x}},{\rm{y}}} \right)}}{{{\rm{\sigma }}\left( {\rm{x}} \right){\rm{\sigma }}\left( {\rm{y}} \right)}}
\)

\(\Rightarrow {\rm{\sigma }}\left( {\rm{x}} \right) = \frac{{{\rm{Cov}}\left( {{\rm{x}},{\rm{y}}} \right)}}{{{\rm{\sigma }}\left( {\rm{y}} \right)\; \times \;{\rm{r}}\left( {{\rm{x}},{\rm{y}}} \right)}}
\)

\(\Rightarrow {\rm{\sigma }}\left( {\rm{x}} \right) = \frac{{27}}{{5\; \times \;0.6{\rm{\;}}}}\)

⇒ σ(x) = 9

Variance of x, σ2(x) = 81

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1767

National Defence Academy (NDA) 2018• Mathematics• Mathematics

The probabilities that a student will solve Question A and Question B are 0.4 and 0.5 respectively. What is the probability that he solves at least one of the two questions? 

  1. AOption (a)

    0.6

  2. BOption (b)

    0.7

  3. COption (c)

    0.8

  4. DOption (d)

    0.9

View answer & explanation

Correct answer

Answer: (b)B.

0.7

Why this answer is correct

Concept:

Formulas used:

  • P(A) = 1 – P(A̅)
  • De Morgan’s law:

\({\rm{P}}\left( {{\rm{\bar A}} \cap {\rm{\bar B}}} \right) = {\rm{P}}(\overline {{\rm{A}} \cup {\rm{B}})}\)

If A and B are two independent events, then:

P(A ∩ B) = P(A) × P(B)

Calculation:

Let,

P(A): probability of solving question A.

P(B): probability of solving question B.

Given: P(A) = 0.4 & P(B) = 0.5

Probability of solving at least one of two questions = P(A ∪ B)

\(= 1 - {\rm{P}}(\overline {{\rm{A}} \cup {\rm{B}})}\)

= 1 – P(A̅ ∩ B̅)      [Using De Morgan’s Law]

= 1 – {P(A̅) × P(B̅)}      [∵ A & B are independent]

= 1 – {[1 – P(A)] × [1 – P(B)]}

= 1 – {(1 – 0.4) × (1 – 0.5)}

= 1 – (0.6 × 0.5)

= 1 – 0.3

= 0.7

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1768

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Let x̅ be the mean of x1, x2, x3, ……xn­ if xi = a + cyi for some constants a and c, then what will be the mean of y1, y2, y3,…, yn?

  1. AOption (a)

    a + c x̅ 

  2. BOption (b)

    \({\rm{a}} - \frac{1}{{\rm{c}}}{\rm{\bar x}}\)

  3. COption (c)

    \(\frac{1}{{\rm{c}}}{\rm{\bar x}} - {\rm{a}}\)

  4. DOption (d)

    \(\frac{{{\rm{\bar x}} - {\rm{a}}}}{{\rm{c}}}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{{{\rm{\bar x}} - {\rm{a}}}}{{\rm{c}}}\)

Why this answer is correct

Concept:

Mean: It is the average of given observation. Let x1, x2, …… , xn be n observations, then

Mean \(= {\rm{\bar X}} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\rm{x}}_{\rm{i}}}}}{{\rm{n}}}\)

Properties:

  • If C is added/subtracted to each observation, then the new mean is also increased/decreased by C.
  • If C is multiplied to each observation, then the new mean is also C times.
  • If each observation is divided by C, then the new mean is old mean divided by C.

Calculation:

Given: \(\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\rm{x}}_{\rm{i}}}}}{{\rm{n}}} = {\rm{\bar x}}\)

Let y̅ be the mean of y1, y2, y3, …, yn

\(\frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} ({\rm{a}} + {\rm{c}}{{\rm{y}}_{\rm{i}}}){\rm{\;}}}}{{\rm{n}}} = {\rm{\bar x}}\)

We know that, if C is subtracted to each observation, then the new mean is also decreased by C.

\(\Rightarrow \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {\rm{c}}{{\rm{y}}_{\rm{i}}}}}{{\rm{n}}} = {\rm{\bar x}} - {\rm{a}}\)

We know that, if each observation is divided by C, then the new mean is old mean divided by C.

\(\Rightarrow \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{n}} {{\rm{y}}_{\rm{i}}}}}{{\rm{n}}} = \frac{{{\rm{\bar x}} - {\rm{a}}}}{{\rm{c}}}\)

\( \Rightarrow {\rm{\bar y}} = \frac{{{\rm{\bar x}} - {\rm{a}}}}{{\rm{c}}}\)

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1769

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Consider the following

Statements:

  1. If the correlation coefficient rxy = 0, then the two lines of regression are parallel to each other.

  2. If the correlation coefficient rxy = + 1, then the two lines of regression are perpendicular to each other.

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (d)D.

Neither 1 nor 2

Why this answer is correct

Concept:

Properties of regression line and correlation coefficient:

  • If the correlation coefficient rxy = 0, then the two lines of regression are perpendicular to each other.
  • If the correlation coefficient rxy = + 1, then the two lines of regression are parallel to each other.

Calculation:

Properties:

  • If the correlation coefficient rxy = 0, then the two lines of regression are perpendicular to each other.
  • If the correlation coefficient rxy = + 1, then the two lines of regression are parallel to each other.

So, statement (1) & statement (2) both are wrong.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1770

National Defence Academy (NDA) 2018• Mathematics• Mathematics

If 4x – 5y + 33 = 0 and 20x – 9y = 107 are two lines of regression, then what are the values of x̅ and y̅ respectively?

  1. AOption (a)

    12 and 18

  2. BOption (b)

    18 and 12

  3. COption (c)

    13 and 17

  4. DOption (d)

    17 and 13

View answer & explanation

Correct answer

Answer: (c)C.

13 and 17

Why this answer is correct

Concept:

Regression lines: Let two variables x and y, if y depends on x, then the result comes in form of simple regression. 

Regression line for simple regression:

y = a + bx

Where a and b are regression parameters.

Properties:

  • The two regression lines (of y on x and x on y) are known to intersect at a specific point (x̅, y̅).

Calculation:

Given: Regression lines,

4x – 5y + 33 = 0      ----(1)

20x – 9y = 107      ----(2)

Now, we know that, intersection point of regression lines gives a point (x̅, y̅).

Multiplying equation (1) with -5 we get,

-20x + 25y – 165 = 0      ----(3)

On adding (2) & (3) we get,

20x – 9y – 107 – 20x + 25y – 165 = 0

⇒ 16y – 272 = 0

⇒ 16y = 272

⇒ y = 17

Put y = 17 in equation (1) we get,

4x – 5(17) + 33 = 0

⇒ 4x = 85 – 33

⇒ 4x = 52

⇒ x = 13

So, x̅ = 13 and y̅ = 17.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1771

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Consider the following

Statements:

  1. Mean is independent of change in scale and change in origin.

  2. Variance is independent of change in scale but not in origin.

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (d)D.

Neither 1 nor 2

Why this answer is correct

Concept:

Properties of mean:

  • If C is added to each observation, then the new mean is also increased by C.
  • If C is multiplied to each observation, then the new mean is also C times.
  • If C is divided to each observation, then the new mean is also old mean divided by C.

Properties of variance:

  • Var(X + C) = Var(X), where C is a constant.
  • Var(CX) = C2 × Var(X), where C is a constant.

Calculation:

Mean:

Change in scale:

We know that, if C is multiplied to each observation, then new mean is also C times old mean.

So, mean is dependent on change in scale.

Change in origin:

We know that, if C is added/subtracted to each observation, then new mean is also increased/decreased by C.

So, mean is dependent on change in origin.

Variance:

Change in scale:

We know that, Var(CX) = C2 × Var(X), where C is a constant.

So, variance is dependent on change in scale.

Change in origin:

We know that, Var(X + C) = Var(X), where C is a constant.

So, variance is independent of change in origin.

Hence, both statement (1) & (2) are incorrect.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1772

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Consider the following

Statements:

  1. The sum of deviations from mean is always zero.

  2. The sum of absolute deviations is minimum when taken around median.

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (c)C.

Both 1 and 2

Why this answer is correct

Concept:

Properties of mean deviation:

  • The sum of deviations from the mean is always zero, i.e., ∑(xi – x̅) = 0.
  • The sum of absolute deviations is minimum when taken around median.

Calculation:

Properties of mean deviation:

  • The sum of deviations from the mean is always zero, i.e., ∑(xi – x̅) = 0.
  • The sum of absolute deviations is minimum when taken around median.

So, statement (1) & statement (2) both are correct.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1773

National Defence Academy (NDA) 2018• Mathematics• Mathematics

What is the median of the numbers 4.6, 0, 9.3, -4.8, 7.6, 2.6, 12.7, 3.5, 8.2, 6.1, 3.9, 5.2?

  1. AOption (a)

    3.8

  2. BOption (b)

    4.9

  3. COption (c)

    5.7

  4. DOption (d)

    6.0

View answer & explanation

Correct answer

Answer: (b)B.

4.9

Why this answer is correct

Concept:

Median: Let a data of n observations is given, then for finding median first arrange the observation either in increasing order or decreasing order.

  • If n is odd, median \(= {\left( {\frac{{\rm{n}}}{2}} \right)^{{\rm{th}}}}\) term.
  • If n is even, median \(= \frac{{{{\left( {\frac{{\rm{n}}}{2}} \right)}^{{\rm{th}}}}{\rm{term\;}} + {{\left( {\frac{{{\rm{n}} }}{2}+1} \right)}^{{\rm{th}}}}{\rm{term}}}}{2}\)

Calculation:

Arranging the given observation in increasing order,

-4.8, 0, 2.6, 3.5, 3.9, 4.6, 5.2, 6.1, 7.6, 8.2, 9.3, 12.7

Number of observation, n = 12

So, we know that if n is even for median then,

\(Median= \frac{{{{\left( {\frac{{\rm{n}}}{2}} \right)}^{{\rm{th}}}}{\rm{term\;}} + {{\left( {\frac{{{\rm{n}} }}{2}+1} \right)}^{{\rm{th}}}}{\rm{term}}}}{2}\)

\( = \frac{{4.6\; +\; 5.2}}{2}\)

= 9.8/2

= 4.9

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1774

National Defence Academy (NDA) 2018• Mathematics• Mathematics

In a test in Mathematics, 20% of the students obtained “first class”. If the data are represented by a Pie-Chart, what is the central angle corresponding to “first class”? 

  1. AOption (a)

    20°

  2. BOption (b)

    36°

  3. COption (c)

    72°

  4. DOption (d)

    144°

View answer & explanation

Correct answer

Answer: (c)C.

72°

Why this answer is correct

Concept:

Formula used:

  • If x% is given as n, then: \({\rm{y\% }} = {\rm{n}} \times \frac{{{\rm{y\% }}}}{{{\rm{x\% }}}}\)

Calculation:

Let the circle be the total number of students given mathematics test.

So, 100% will give 360°.

∴ 20% will be \(360^\circ \times \frac{{20{\rm{\% }}}}{{100{\rm{\% }}}}{\rm{\;}}\)

= 72°

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1775

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Two fair dice are rolled. What is the probability of getting a sum of 7?

  1. AOption (a)

    1/36

  2. BOption (b)

    1/6

  3. COption (c)

    7/12

  4. DOption (d)

    5/12

View answer & explanation

Correct answer

Answer: (b)B.

1/6

Why this answer is correct

Concept:

Probability: It is a measure of likelihood of an event to occur.

Probability of an event to occur, P(E) \(= \frac{{{\rm{number\;of\;favourable\;outcomes}}}}{{{\rm{number\;of\;total\;outcomes}}}}\)

Formula used:

  • If a dice is rolled n times then, number of sample space n(S) = 6n

Calculation:

Let,

S: Event of rolling 2 dice.

A: Event that the sum of the numbers appeared on both dice to be 7.

Number of total outcomes = n(S)

= 62

= 36

Now, A = {(1,6), (2,5), (3,4), (4,3), (5,2), (6,1)}

Number of favourable outcomes = n(A)

= 6

Probability of getting sum of 7, P(A) \(= \frac{{{\rm{number\;of\;favourable\;outcomes}}}}{{{\rm{number\;of\;total\;outcomes}}}}\)

= 6/36

= 1/6

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1776

National Defence Academy (NDA) 2018• Mathematics• Mathematics

If A and B are two events such that 2P(A) = 3P(B), where 0 < P(A) < P(B) < 1, then which one of the following is correct?

  1. AOption (a)

    P(A|B) < P(B|A) < P(A ∩ B)

  2. BOption (b)

    P(A ∩ B) < P(B|A) < P(A|B)

  3. COption (c)

    P(B|A) < P(A|B) < P(A ∩ B)

  4. DOption (d)

    P(A ∩ B) < P(A|B) < P(B|A)

View answer & explanation

Correct answer

Answer: (b)B.

P(A ∩ B) < P(B|A) < P(A|B)

Why this answer is correct

Concept:

Formulas used:

  • Conditional probability:

\({\rm{P}}({\rm{A}}|{\rm{B}}) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}}\)

Calculation:

Given: 2P(A) = 3P(B) & 0 < P(A) < P(B) < 1

Dividing both sides by P(A ∩B) to both sides we get,

\( \Rightarrow \frac{{2{\rm{P}}\left( {\rm{A}} \right)}}{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}} = \frac{{3{\rm{P}}\left( {\rm{B}} \right)}}{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}\)

Taking reciprocal of both sides,

\(\Rightarrow \frac{1}{2} \times \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{A}} \right)}} = \frac{1}{3} \times \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}}\)

⇒ (1/2) × P(B|A) = (1/3) × P(A|B)

⇒ P(B|A) = (2/3) P(A|B)

So, P(B|A) < P(A|B)        ----(1)

From conditional probability,

\({\rm{P}}\left( {{\rm{B|A}}} \right) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{A}} \right)}}\)

Given, P(A) < 1

So, P(B|A) > P(A ∩ B)        ----(2)

\({\rm{P}}({\rm{A}}|{\rm{B}}) = \frac{{{\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}}\)

Given, P(B) < 1

So, P(A|B) > P(A ∩ B)        ----(3)

From (1), (2) & (3)

P(A ∩ B) < P(B|A) < P(A|B)

Hence, option (2) is correct.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1777

National Defence Academy (NDA) 2018• Mathematics• Mathematics

A box has ten chits numbered 0, 1, 2, 3, ……., 9. First, one chit is drawn at random and kept aside. From the remaining, a second chit is drawn at random. What is the probability that the second chit drawn is “9”?

  1. AOption (a)

    1/10

  2. BOption (b)

    1/9

  3. COption (c)

    1/90

  4. DOption (d)

    None of the above

View answer & explanation

Correct answer

Answer: (a)A.

1/10

Why this answer is correct

Concept:

Probability: It is a measure of likelihood of an event to occur.

Probability of an event to occur, P(E) \(= \frac{{{\rm{number\;of\;favourable\;outcomes}}}}{{{\rm{number\;of\;total\;outcomes}}}}\)

Note:

  • Combination is the number of ways in which we can select a group of objects from a set. For example, selecting r things from set of n things is given as nCr.

Calculation:

Given: A box has 10 chits numbered {0, 1, 2, …,9}.

Let,

S: Event of drawing two chits one by one.

A: Event of second chit drawn is numbered 9.

So, number of total outcomes = n(S)

= 10C1 × 9C1

= 10 × 9

= 90

Number of favourable outcomes = n(A) = (0,9) (1,9).........(8,9)

= 9

Probability of second chit drawn to be 9 = \(\frac{{{\rm{n}}\left( {\rm{A}} \right)}}{{{\rm{n}}\left( {\rm{S}} \right)}}\)

= 9/90 = 1/10

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1778

National Defence Academy (NDA) 2018• Mathematics• Mathematics

Consider the following in respect of two events A and B:

  1. P(A occurs but not B) = P(A) – P(B) if B ⊂ A

  2. P(A alone or B alone occurs) = P(A) + P(B) – P(A ∩ B)

  3. P(A ∪ B) = P(A) + P(B) if A and B are mutually exclusive

Which of the above is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    1 and 3 only

  3. COption (c)

    2 and 3 only

  4. DOption (d)

    1 and 2 only

View answer & explanation

Correct answer

Answer: (b)B.

1 and 3 only

Why this answer is correct

Concept:

Formulas used:

Let A and B be two subsets of X.

  • P(A – B) = P(A) – P(B) if B ⊂ A
  • P( A alone) = P(A) – P(A ∩ B)
  • P(A ∪ B) = P(A) + P(B) – P(A ∩ B)
  • If A and B are mutually exclusive, then P(A ∩ B) = 0

Calculation:

P(A occurs but not B) = P(A – B)

= P(A) – P(B)

So, statement (1) is correct.

P(A alone or B alone occurs) = P(A) – P(A ∩ B) + P(B) – P(A ∩ B)

= P(A) + P(B) – 2P(A ∩ B)

So, statement (2) is incorrect.

P(A ∪ B) = P(A) + P(B) – P(A ∩ B)

Since, A and B are mutually exclusive, P(A ∩ B) = 0

P(A ∪ B) = P(A) + P(B)

So, statement (3) is correct.

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1779

National Defence Academy (NDA) 2018• Mathematics• Mathematics

The standard deviation σ of the first N natural numbers can be obtained using which one of the following formulae?

  1. AOption (a)

    \({\rm{\sigma }} = \frac{{{{\rm{N}}^2} - 1}}{{12}}\)

  2. BOption (b)

    \({\rm{\sigma }} = \sqrt {\frac{{{{\rm{N}}^2} - 1}}{{12}}} \)

  3. COption (c)

    \({\rm{\sigma }} = \sqrt {\frac{{{\rm{N}} - 1}}{{12}}}\)

  4. DOption (d)

    \({\rm{\sigma }} = \sqrt {\frac{{{{\rm{N}}^2} - 1}}{{6{\rm{N}}}}}\)

View answer & explanation

Correct answer

Answer: (b)B.

\({\rm{\sigma }} = \sqrt {\frac{{{{\rm{N}}^2} - 1}}{{12}}} \)

Why this answer is correct

Concept:

Formulas used:

  • Variance, \({{\rm{\sigma }}^2} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {{\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)}^2}}}{{\rm{N}}}\)
  • Standard deviation = σ
  • Summation of square of first N natural number,  \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{N}} {\rm{x}}_{\rm{i}}^2 = \frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)\left( {2{\rm{N}} + 1} \right)}}{6}\)
  • Summation of first N natural number, \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{N}} {{\rm{x}}_{\rm{i}}} = 1 + 2 + 3 + \ldots + {\rm{N}} = \frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)}}{2}\)
  • \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{N}} 1 = {\rm{N}}\)

Properties of summation:

  • \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} \left( {{{\rm{a}}_{\rm{i}}} + {{\rm{b}}_{\rm{i}}}} \right) = \mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} \left( {{{\rm{a}}_{\rm{i}}}} \right) + \mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} \left( {{{\rm{b}}_{\rm{i}}}} \right)\)
  • \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} \left( {{\rm{c}}{{\rm{a}}_{\rm{i}}}} \right) = {\rm{c}}\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} \left( {{{\rm{a}}_{\rm{i}}}} \right)\)
  • \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} {\left( {{{\rm{a}}_{\rm{i}}} + {{\rm{b}}_{\rm{i}}}} \right)^2} = \mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} {\left( {{{\rm{a}}_{\rm{i}}}} \right)^2} + \mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} {\left( {{{\rm{b}}_{\rm{i}}}} \right)^2}\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{n}} \left( {{{\rm{a}}_{\rm{i}}}{{\rm{b}}_{\rm{i}}}} \right)\)

Calculation:

To find: σ of first N natural number

Let xi = 1, 2, 3, …,n for i = 1, 2, 3, …,n.

Let x̅ be mean of xi for i = 1, 2, 3, …,n.

Variance, \({{\rm{\sigma }}^2} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {{\left( {{{\rm{x}}_{\rm{i}}} - {\rm{\bar x}}} \right)}^2}}}{{\rm{N}}}\)

\( \Rightarrow {{\rm{\sigma }}^2} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} ({\rm{x}}_{\rm{i}}^2 + {{{\rm{\bar x}}}^2} - 2{{\rm{x}}_{\rm{i}}}{\rm{\bar x}})}}{{\rm{N}}}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {\rm{x}}_{\rm{i}}^2{\rm{\;}} + {\rm{\;}}\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {{{\rm{\bar x}}}^2}{\rm{\;}} - {\rm{\;}}\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} 2{{\rm{x}}_{\rm{i}}}{\rm{\bar x}}}}{{\rm{N}}}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {\rm{x}}_{\rm{i}}^2{\rm{\;}} + {\rm{\;}}{{{\rm{\bar x}}}^2}\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} 1 - {\rm{\;}}2{\rm{\bar x}}\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {{\rm{x}}_{\rm{i}}}}}{{\rm{N}}}\)

We know that,

Summation of square of first N natural number, \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{N}} {\rm{x}}_{\rm{i}}^2 = \frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)\left( {2{\rm{N}} + 1} \right)}}{6}\)

Summation of first N natural number, \(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{N}} {{\rm{x}}_{\rm{i}}} = 1 + 2 + 3 + \ldots + {\rm{N}} = \frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)}}{2}\)

Mean of first N natural number, \({\rm{\bar x}} = \frac{{1 + 2 + 3 + \ldots + {\rm{N}}}}{{\rm{N}}}\)

\(\Rightarrow {\rm{\bar x}} = \frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)}}{{2{\rm{N}}}}\)

\(\Rightarrow {\rm{\bar x}} = \frac{{{\rm{N}} + 1}}{2}\)

\(\mathop \sum \limits_{{\rm{i}} = 1}^{\rm{N}} 1 = {\rm{N}}\)

\({{\rm{\sigma }}^2} = \frac{{\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {\rm{x}}_{\rm{i}}^2{\rm{\;}} + {\rm{\;}}{{{\rm{\bar x}}}^2}\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} 1 - {\rm{\;}}2{\rm{\bar x}}\mathop \sum \nolimits_{{\rm{i}} = 1}^{\rm{N}} {{\rm{x}}_{\rm{i}}}}}{{\rm{N}}}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)\left( {2{\rm{N}} + 1} \right)}}{6}{\rm{\;}} + {\rm{\;}}{{\left( {\frac{{{\rm{N}} + 1}}{2}} \right)}^2} \times {\rm{N\;}} - {\rm{\;}}2 \times \left( {\frac{{{\rm{N}} + 1}}{2}} \right) \times \frac{{{\rm{N}}\left( {{\rm{N}} + 1} \right)}}{2}}}{{\rm{N}}}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\left( {{\rm{N}} + 1} \right)\left( {2{\rm{N}} + 1} \right)}}{6} + {\left( {\frac{{{\rm{N}} + 1}}{2}} \right)^2} - {\left( {\frac{{{\rm{N}} + 1}}{4}} \right)^2}\)

\( \Rightarrow {{\rm{\sigma }}^2} = \frac{{\left( {{\rm{N}} + 1} \right)\left( {2{\rm{N}} + 1} \right)}}{6} - {\left( {\frac{{{\rm{N}} + 1}}{4}} \right)^2}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\left( {{\rm{N}} + 1} \right)\left\{ {2 \times \left( {2{\rm{N}} + 1} \right) - 3 \times \left( {{\rm{N}} + 1} \right)} \right\}}}{{12}}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\left( {{\rm{N}} + 1} \right)\left\{ {4{\rm{N}} + 2 - 3{\rm{N}} - 3} \right\}}}{{12}}\)

\(\Rightarrow {{\rm{\sigma }}^2} = \frac{{\left( {{\rm{N}} + 1} \right)\left( {{\rm{N}} - 1} \right)}}{{12}}\)

\( \Rightarrow {{\rm{\sigma }}^2} = \frac{{{{\rm{N}}^2} - 1}}{{12}}\)

\(\Rightarrow {\rm{\sigma }} = \sqrt {\frac{{{{\rm{N}}^2} - 1}}{{12}}} \)

NDA Mathematics 22 April 2018 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1780

National Defence Academy (NDA) 2018• Mathematics• Mathematics

The coordinates of the vertices P, Q and R of a triangle PQR are (1, -1, 1) (3, -2, 2) and (0, 2, 6) respectively. If ∠RQP = θ, Then what is ∠PRQ equal to?

  1. AOption (a)

    30° + θ 

  2. BOption (b)

    45° - θ

  3. COption (c)

    60° - θ

  4. DOption (d)

    90° - θ

View answer & explanation

Correct answer

Answer: (d)D.

90° - θ

Why this answer is correct

Concept:

  1. If two vector a and b then \({\rm{a}}.{\rm{b}} = \left| {{\rm{\vec a}}} \right|{\rm{\;}}\left| {{\rm{\vec b}}} \right|\cos {\rm{\theta }}\)

 Where θ is angle between vector a and b.

  1. Vector A = \({\rm{x\hat i}} + {\rm{y\hat j}} + {\rm{z\hat k}}\) then |A| = \(\sqrt {{{\rm{x}}^2} + {{\rm{y}}^2} + {{\rm{z}}^2}} \)

 

Calculation:

Given that,

P (1, -1, 1), Q (3, -2, 2) and R (0, 2, 6)

So from diagram

\(\begin{array}{l} \overrightarrow {{\rm{PR}}} = {\rm{\;}} - 1{\rm{\hat i}} + 3{\rm{\hat j}} + 5\hat k\\ \overrightarrow {{\rm{PQ}}} = {\rm{\;}}2{\rm{\hat i}} - {\rm{\hat j}} + \hat k \end{array}\)

Now, Angle between two vectors are

\(\begin{array}{l} \overrightarrow {{\rm{PR}}} {\rm{\;}}.\overrightarrow {{\rm{PQ}}} = \left| {\overrightarrow {{\rm{PQ}}} } \right|{\rm{\;}}\left| {\overrightarrow {{\rm{PR}}} } \right|\cos {\rm{\theta \;}}\\ \Rightarrow \overrightarrow {{\rm{PR}}} {\rm{\;}}.\overrightarrow {{\rm{PQ}}} = \left( { - 1{\rm{\hat i}} + 3{\rm{\hat j}} + 5{\rm{\hat k}}} \right).\left( {2{\rm{\hat i}} - {\rm{\hat j}} + {\rm{\hat k}}} \right)\\ \Rightarrow \overrightarrow {{\rm{PR}}} {\rm{\;}}.\overrightarrow {{\rm{PQ}}} = {\rm{\;}} - 2 - 3 + 5 = 0 \end{array}\)

So,

\(\left| {\overrightarrow {{\rm{PQ}}} } \right|{\rm{\;}}\left| {\overrightarrow {{\rm{PR}}} } \right|\cos \emptyset = \overrightarrow {{\rm{PR}}} {\rm{\;}}.\overrightarrow {{\rm{PQ}}} = 0\)

∴ cos ∅ = 0

⇒ ∅ = 90°

As we know,

Sum of all internal angle of tringle equal to 180

So from the diagram,

⇒ ∅ + θ + x = 180°

Put ∅ = 90°

⇒ ∠PRQ = x = 90 - θ

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.